x+3y-24=0
change each equation to slope intercept form and state the slope and y intercept

Answer by nerdybill(7384)

You are trying to manipulate the equation into the "slope-intercept" form of:
y = mx + b
where
m is slope
b is y-intercept
.
Starting with:
x+3y-24=0
subtract x from both sides:
3y-24 = -x
add 24 to both sides:
3y = -x + 24
divide both sides by 3:
y = (-1/3)x + 8
or
y = -x/3 + 8